OpenStudy (anonymous):

Can someone take me step by step on how evaluating this expression and finding variable? 7/x^2-5x+4 x=-1

OpenStudy (anonymous):

replace \(x\) by \((-1)\) use the parentheses, since it is a negative number

OpenStudy (anonymous):

you should get \[\frac{7}{(-1)^2-5\times (-1)+4}\]

OpenStudy (anonymous):

Yes, this is what I have so far.

OpenStudy (anonymous):

then you compute the number \[\frac{7}{(-1)^2-5\times (-1)+4}\] \[=\frac{7}{1+5+4}=\frac{7}{10}\]

OpenStudy (anonymous):

And that is it?

OpenStudy (anonymous):

wow, easier than I thought